OpenStudy (anonymous):

EASY QUESTION! Hope uses the equation C = 3 h + 9 to find the total cost, C , in dollars, of renting a bike for h hours. If Hope does not spend more than $30, what is the maximum number of hours she can rent the bike? A 18 B 13 C 10 D 7

OpenStudy (anonymous):

13 or 10

OpenStudy (ivettef365):

use equation, so you have 30 < 3h + 9 solve for h

OpenStudy (ivettef365):

how did you get 27?? you have c = 3(7) + 9 c = 21 +9 c = 30 so that is correct (7)

OpenStudy (ivettef365):

yw :)